OpenStudy (anonymous):

Factor: x2 + 11x + 10 (x + 10)(x +1) (x + 11)(x + 10) (x + 2)(x + 5) prime

OpenStudy (anonymous):

(x + 10)(x +1)

OpenStudy (anonymous):

(x+11) (x-1)

OpenStudy (anonymous):

cocotsoi this is the correct check ur answer

OpenStudy (anonymous):

(x + 10)(x +1) = x^2 +10x +x+10 = x^2 +11x+10